If a woman works for 3 1/2 hours per day, she can finish a piece of work in 10 days. How many hours per day should she work to finish the same piece of work in 7 days?
Answer by Theo(13342)

if she works 3.5 hour a day and it takes her 10 days to finish a piece of work, then it takes her 3.5 * 10 days = 35 hours to finish the piece of work.

the amount of work required is 35 hours to finish the job.

this remains the same.

let x = the number of hours it takes her to finish the finish the job in 7 days.

you get x * 7 = 35

solve for x to get:

x = 5

it would take her 5 hours per day to finish the job in 7 days.

5 * 7 = 35
3.5 * 10 = 35